OpenStudy (anonymous):

7th grade math help please

OpenStudy (anonymous):

Simplify. −1w + 4w4 + y4 – 3y4 – 2w – 3w4 A. −2w4 – 2y4 B. −w4y4 – 3w C. w4 – 3w – 2y4 D. 12w4 + 2w – 3y4

OpenStudy (anonymous):

leave then come back in to see it better:)

OpenStudy (anonymous):

are you allowed to use a actual good calculator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

take all the corresponding letters and add them together like ex:4w+3w

OpenStudy (anonymous):

it would be 7w but thats a example

OpenStudy (anonymous):

i had got -23 but idk if it is right???:(

OpenStudy (anonymous):

if im thinking right thats what you would do

OpenStudy (anonymous):

can i get more help please on the same problem please:)

OpenStudy (texaschic101):

I see it just fine. you have to comine like terms -1w + 4w^4 + y^4 - 3y^4 - 2w - 3w^4 I am going to seperate this so you can see the like terms -1w - 2w = -3w 4w^4 - 3w^4 = w^4 y^4 - 3y^4 = -2y^4 so you are left with : w^4 - 3w - 2y^4

OpenStudy (anonymous):

so C

OpenStudy (texaschic101):

yep...C

OpenStudy (anonymous):

there ya go got a math genious in here :D good job

OpenStudy (anonymous):

Thank You soooooooooooooooooooooooooo much:) I have more ok

OpenStudy (texaschic101):

I don't know about a genius...lol...but I try :)

OpenStudy (texaschic101):

what else ya got ?

OpenStudy (anonymous):

You r an genius

OpenStudy (anonymous):

The growth of one child is represented by the expression 5x – 1. The growth of a second child is represented by the expression 2x + 3. Complete the expression to represent the total growth of both children. ______x + _________

OpenStudy (anonymous):

leave then come back in to see it:)

OpenStudy (texaschic101):

hold on...got to refresh

OpenStudy (anonymous):

*better

OpenStudy (anonymous):

@texaschic101 after you help me i have someone that needs help ok:)

OpenStudy (texaschic101):

better...pain in the butt black diamonds...lol 5x - 1 + 2x + 3 = 7x + 2

OpenStudy (anonymous):

April adds the expressions 9x + 7 and 5x – 3. Evan adds the expressions 14x + 2 and 4x + 3. How much greater is Evan’s sum than April’s sum? __________x + _______

OpenStudy (anonymous):

I do not get how to do thisone @texaschic101

OpenStudy (anonymous):

can you help me with this one this is the last one please

OpenStudy (anonymous):

and after you help me @Rainbow_dashie needs your help:) If you do not mind:)

OpenStudy (rainbow_dashie):

Which equations model this problem? A recipe that makes 1 serving calls for 0.25 L milk. How many servings can Rohit make if he uses 2 L of milk? Use s for the number of servings. Choose all answers that are correct. A. B. C. D.

OpenStudy (rainbow_dashie):

hold on there are files

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@texaschic101 @texaschic101 @texaschic101 @texaschic101 @texaschic101

OpenStudy (anonymous):

@texaschic101 :):):):):)

OpenStudy (anonymous):

omg i only need help witht his last question:(

OpenStudy (anonymous):

What unit an lesson is this

OpenStudy (anonymous):

that you are doing @Rainbow_dashie

OpenStudy (texaschic101):

9x + 7 + 5x - 3 = 14x + 4 14x + 2 + 4x + 3 = 18x + 5 difference is : (18x + 5) - (14x + 4) -- distribute through the parenthesis 18x + 5 - 14x - 4 = 4x + 1 ================= 1/0.25 = x/2 --- 1 serving to 0.25 liters = x servings to 2 liters cross multiply 0.25x = 2 -- divide by 0.25 x = 8 servings

OpenStudy (rainbow_dashie):

ok

OpenStudy (texaschic101):

@cjackson2002 ....what happened ?

OpenStudy (anonymous):

do you remember the first one that you put out here put it in this box ok i have the right answer:)

OpenStudy (anonymous):

@Rainbow_dashie

OpenStudy (texaschic101):

Is it the one with April and Evan ??

OpenStudy (texaschic101):

or the simplifying one ??

OpenStudy (texaschic101):

which one did you miss ??

OpenStudy (texaschic101):

ok...well tag me if you need help